OpenStudy (anonymous):

If the linear equation 4x = 10 - 5y is written in standard form, what is the value of B? Answer

OpenStudy (rulnick):

4x+5y=10 B=5

OpenStudy (rulnick):

... Ax+By=C being "standard form."

OpenStudy (anonymous):

if standard form is y = mx + b, then b is 2. the equation is y = 4/5(x) + 2